### v1

# Stepwise Reasoning Error Disruption Attack of LLMs

Abstract
Large language models (LLMs) have made remarkable strides in complex reasoning tasks, but their safety and robustness in reasoning processes remain underexplored. Existing attacks on LLM reasoning are constrained by specific settings or lack of imperceptibility, limiting their feasibility and generalizability. To address these challenges, we propose the Stepwise rEasoning Error Disruption (SEED) attack, which subtly injects errors into prior reasoning steps to mislead the model into producing incorrect subsequent reasoning and final answers. Unlike previous methods, SEED is compatible with zero-shot and few-shot settings, maintains the natural reasoning flow, and ensures covert execution without modifying the instruction. Extensive experiments on four datasets across four different models demonstrate SEED’s effectiveness, revealing the vulnerabilities of LLMs to disruptions in reasoning processes. These findings underscore the need for greater attention to the robustness of LLM reasoning to ensure safety in practical applications.
[…]
Stepwise Reasoning Error Disruption Attack of LLMs
Jingyu Peng‡§111 Both authors contributed equally to this research., Maolin Wang§111 Both authors contributed equally to this research., Xiangyu Zhao§ 222 Corresponding authors. , Kai Zhang, Wanyu Wang§, Pengyue Jia§, Qidong Liu§, Ruocheng Guo, Qi Liu 222 Corresponding authors. University of Science and Technology of China, § City University of Hong Kong, ByteDance Research jypeng28@mail.ustc.edu.cn morin.wang@my.cityu.edu.hk
1 Introduction
Large language models (LLMs) have remarkably improved complex tasks by adopting various enhanced reasoning approaches Besta et al. (2024); Yang et al. (2024); Yao et al. (2024). These approaches have boosted their performance and drawn attention to the trustworthiness of the reasoning processes, including faithfulness Lanham et al. (2023); Turpin et al. (2024), fairness Shaikh et al. (2023), and safety Xu et al. (2024).
Previous work has exposed significant LLM vulnerabilities in simple tasks such as classification and generation Wang et al. (2024a); Zhao et al. (2023); Xu et al. . However, their susceptibility to attacks during the complex reasoning processes - where the stakes are often higher and the consequences are more severe in some critical area - remains largely unexplored. Recent advances in reasoning methods have introduced effective mechanisms, enabling models to iteratively build upon prior steps: from basic chain-of-thought approaches Wei et al. (2022); Kojima et al. (2022) to advanced reflection mechanisms  Madaan et al. (2024); Zhao et al. (2024) and multi-step reasoning Wu et al. (2024); Wang et al. (2024b). This critical dependence on step-wise reasoning introduces a new type of vulnerability in LLMs, where manipulation of initial reasoning steps can propagate errors, causing cascading failures throughout the reasoning chain.
Exploiting such vulnerability in LLMs introduces two fundamental challenges: feasibility and imperceptibility. Technically, unlike traditional adversarial attack methods, which often leverage internal information of models such as gradients and logits, state-of-the-art LLMs are now primarily deployed as proprietary APIs Achiam et al. (2023); Team et al. (2023). Therefore, only prompt-based attacks are feasible, where adversaries have to operate through input manipulation. While existing attempts to compromise LLM reasoning Xu et al. (2024); Xiang et al. (2024) have demonstrated success in specific scenarios, they still face severe limitations in practice. A key challenge in attack design is crafting imperceptible attacks. While obvious manipulations like changing final answers or inserting irrelevant steps are easily detected, modifying the reasoning output while maintaining narrative coherence is rather challenging. Existing approaches struggle to achieve this balance between attack effectiveness and stealth in complex reasoning tasks.
Among the most relevant approaches, Xiang et al. (2024) employs misleading demonstrations to induce errors in LLMs. However, these methods are limited to in-context learning scenarios, requiring demonstrations as input, which limits their generalizability to the zero-shot settings. Furthermore, their strategy introduces an additional step that modifies the final answer, making it quite easy to identify. Another related approach, the preemptive answer “attack” Xu et al. (2024) alters the reasoning paradigm of the model by producing conclusions before deriving reasoning steps. Despite its novelty, this approach often generates easily identifiable outputs, reducing its imperceptibility and effectiveness in practice. These limitations are further evidenced by our experimental results in Section 3.2.
To address these two limitations, we propose the Stepwise rEasoning Error Disruption (SEED) attack. First, SEED addresses the feasibility challenge by leveraging LLMs’ reliance on step-by-step reasoning. Instead of depending on demonstrations or backpropagated gradients, SEED strategically introduces subtle errors into the early reasoning steps. This approach achieves high success rates across a wide range of scenarios without the need for task-specific training or examples, proving its effectiveness within the constraints of proprietary API-based LLM deployments in both zero-shot and few-shot settings. Second, SEED overcomes the challenge of imperceptibility by maintaining the original prompt structure while subtly manipulating the reasoning process. The carefully introduced errors seamlessly integrate into the reasoning flow, naturally propagating through the reasoning chain to produce incorrect yet plausible-looking outcomes. This ensures that the disruptions remain covert, avoiding detection while preserving the model’s perceived trustworthiness. This novel approach not only addresses the identified limitations but also introduces a fresh perspective on how reasoning vulnerabilities in LLMs can be exploited.
Our contributions can be summarized as follows:
  • We define the task of disrupting the step-by-step reasoning process of LLMs and introduce SEED, a versatile and effective attack method that is both highly efficient in execution and challenging in detection.
  • We demonstrate the effectiveness and stealth of SEED across four representative LLMs on four datasets with different characteristics, which include diverse and challenging reasoning tasks presented in two different formats.
  • We naturally validate the vulnerability of LLMs to adversarially injected prior reasoning steps by designing SEED, which effectively exploits these weaknesses.
Figure 1: An example demonstrating the definition of a step-by-step reasoning task for an LLMs.
2 Method
In this section, we first provide an explicit definition of attacks that target the step-by-step reasoning process of LLMs. Following that, we introduce our two implementations of the proposed SEED attack.
Figure 2: An example of the SEED-S attack on a math problem. The red font highlights the generated misleading content where subtle calculation errors are introduced while maintaining the reasoning coherence.
2.1 Problem Formulation
We first present a formal definition of a step-by-step reasoning task of LLMs as shown in Fig 1. For a given problem $p$, we define the query to the LLM, denoted as $q$, as follows:
$q=[I_{solve}\,||\,D\,||\,p],$
where $D=[d_{1},\dots,d_{K}]$ and $d_{k}$ represents the $k$-th demonstration in few-shot setting. Each demonstration $d_{k}$ is structured as $[p_{k},[r_{k}^{1},\dots,r_{k}^{T}],a_{k}]$, with $r_{k}^{t}$ being the $t$-th step in the reasoning process for the problem $p_{k}$, and $a_{k}$ representing the final answer. If $K=0$, the setting is reduced to a zero-shot scenario from few-shot.
Given $q$ as input, the corresponding output $o$ of the LLM is expressed as:
$o=LLM(q)=[R\,||\,a],$
where $R_{i}=[r^{1},\dots,r^{T}]$ is the reasoning process. Attacks targeting the reasoning process of LLMs focus on altering $o$ and its corresponding $a$ by modifying $q$ into $q^{\prime}$, which can be formulated as:
$\mathop{\arg\max}\limits_{q^{\prime}}LLM_{a^{\prime}}(q^{\prime})$ (1)
[…]
where $LLM_{a^{\prime}}$ represents the probability of the output answer being equal to $a^{\prime}$ and $\text{diff}(\cdot)$ represents the difference in terms of the narrative structure and semantic similarity.
Figure 3: An example of the SEED-P attack on a math problem. The red font highlights the generated misleading content, while the gray font highlights the optional part for problems with answer choices.
2.2 Overview of Stepwise Reasoning Error Disruption Attack
Due to certain observations (as detailed in Section 3.2), modifications to $I_{solve}$ appear to be easily detectable, which could be partially explained by the sensitivity of the model to perturbations in problem-solving inputs. Similarly, changes to $p$ seem to be detectable by prompting the LLM to repeat the problem, potentially leveraging its tendency toward consistent reasoning in generating responses. Meanwhile, modification on demonstrations is not supported under zero-shot setting. Therefore, SEED attack performs the attack by adding misleading steps $R_{att}=[r_{att}^{1},\dots,r_{att}^{T_{att}}]$ and eliciting the LLM to output the subsequent reasoning steps $R^{\prime}=[r^{\prime 1},\dots,r^{\prime T^{\prime}}]$ and the final answer $a^{\prime}$ based on $R^{\prime}$:
$o^{\prime}=R^{\prime}||a^{\prime}=LLM([I_{solve}||D||p||R_{att}]).$
[…]
$\mathop{\arg\max}\limits_{R_{att}}\;LLM_{a^{\prime}}(I_{solve}||D||p||R_{att})$ (2)
[…]
It’s worth noting that, as we take some reasoning steps $R_{att}$ as input, we will display $[R_{att}||R^{\prime}]$ for the victim user to maintain the integrity of reasoning process. Therefore, the constraint $\text{diff}(R,R^{\prime})$ is converted to $\text{diff}(R,[R_{att}||R^{\prime}])$.
Besides, we assume that the reasoning steps are continuous, with each step depending on the previous ones. Therefore, we can get:
$\text{diff}(R,[R_{att}||R^{\prime}]\propto\text{diff}(R[:T_{att}],R_{att}),$
with the constraint that $T_{att}+T^{\prime}=T$. In practice, as the number of reasoning steps $T$ varies, we introduce $\sigma=\frac{T_{att}}{T}$ as a hyperparameter to control the $T_{att}$. To generate $R_{att}$ that both closely resembles $R[:T_{att}]$ and effectively misleads the LLM into providing an incorrect answer, we developed two LLM-based implementations. In the following two subsections, we propose two implementations of SEED attack: SEED-S (Step Modification) and SEED-P (Problem Modification). SEED-S directly modifies the final step of the reasoning process, while SEED-P generates a modified problem that leads to the desired incorrect answer.
2.3 SEED-S: SEED Attack by Step Modification
As shown in Figure 2, one intuitive and straightforward approach is to modify the final step of $R[:T_{att}]$ with the help of an assistant LLM:
$r_{mod}$ $=LLM_{assist}(I_{mod}||p||R^{\prime}[T_{att}])$ (3)
$R_{att}$ $=R[:T_{att}-1]||r_{mod}),$
where $r_{mod}$ refers to the modified reasoning step and $I_{mod}$ refers to the instruction given to the LLM to modify the reasoning step in a way that leads to an incorrect answer. It is important to note that we instruct the LLM to only modify certain digits or words related to the final answer, rather than regenerate an entirely different step, ensuring that the similarity and length constraint is still met.
However, this naive implementation has some drawbacks. Firstly, it has been observed that LLMs tend to pay more attention to the head and tail parts of the input. As a result, LLMs are more likely to notice inconsistencies between the last two steps. Secondly, the final result $a^{\prime}$ is easily noticeable, for example, if there is a significant discrepancy in magnitude compared to $a$, or if the final answer in a multiple-choice question does not appear among the available options. In such cases, the LLM is likely to recognize and correct the error.
Table 1: A comparison of the proportions of solutions generated by BadChain Xiang et al. (2024), UPA and MPA Xu et al. (2024), and SEED (SEED-S and SEED-P) that were detected by GPT-4o as originating from prompts containing attacks. The average improvement is determined by calculating the average decline in the detection rate of SEED compared to Xu et al. (2024).Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Results demonstrate that SEED methods consistently achieve substantially lower detection rates across all model architectures and settings, with SEED-P showing particularly strong stealth capabilities while maintaining attack effectiveness.
MATH GSM8K
Setting BadChain UPA MPA SEED-S SEED-P Avg. Impr. BadChain UPA MPA SEED-S SEED-P Avg. Impr.
Llama Z_S 0.998 0.382 0.440 0.170 0.252 48.7% 1.000 0.442 0.526 0.088 0.204 69.8%
F_S 1.000 0.260 0.438 0.150 0.208 48.7% 0.998 0.226 0.384 0.066 0.146 65.2%
Qwen Z_S 0.998 0.336 0.325 0.053 0.077 80.3% 0.994 0.484 0.407 0.039 0.166 77.0%
F_S 0.996 0.352 0.382 0.026 0.091 84.1% 0.996 0.439 0.497 0.042 0.162 78.2%
Mistral Z_S 0.998 0.526 0.546 0.219 0.382 43.9% 1.000 0.496 0.494 0.106 0.292 59.8%
F_S 1.000 0.537 0.478 0.212 0.421 37.6% 0.996 0.468 0.408 0.150 0.334 44.7%
GPT4-o Z_S 1.000 0.439 0.353 0.032 0.052 89.4% 1.000 0.502 0.572 0.008 0.042 95.3%
F_S 0.996 0.36 0.362 0.026 0.026 92.8% 0.998 0.426 0.406 0.014 0.022 95.7%
2.4 SEED-P: SEED Attack by Problem Modification.
To solve the limitation of SEED-S due to LLMs’ heightened attention to sequence endings and potential magnitude discrepancies in final answers, we propose a more meticulously designed implementation involving modifying the raw problem, as illustrated in Figure 3. The process begins by prompting the assistant LLM to solve the original problem and obtain the raw answer. With knowledge of this answer, the LLM is more likely to generate a modified problem that is both similar to the original and aligned with its corresponding answer. The whole process can be expressed as:
$p_{mod}||R_{mod}||a_{mod}=LLM_{assist}(p,a).$
By providing more fluent reasoning steps $R_{att}=R_{mod}[:T_{att}]$, the target LLM becomes more susceptible to being misled, ultimately producing incorrect reasoning steps and an incorrect answer.
For reasoning tasks with answer choices, the LLM is first instructed to select an answer choice, and then generate a problem based on the chosen answer. This ensures that the generated question aligns with the provided answer choices, maintaining the necessary consistency for successful attack.
To further enhance the attack’s effectiveness, inspired by Xu et al. (2024), we prepend the corresponding incorrect answer to $R_{att}$. Finally, the modified output of the target LLM is obtained by feeding the modified problem’s incorrect answer and partial reasoning steps into it:
$q^{\prime}=LLM(D_{i}||I||p_{i}||a^{\prime}||R_{att}).$
Since we prepend $a^{\prime}$ to $R_{att}$, the proportion of $a^{\prime}$ relative to the entire input $q^{\prime}$ is minimal, and its position is central. Thus, we assume that its impact on the length of $R^{\prime}$ and the similarity between model outputs $R^{\prime}$ and $R[T_{att}:]$ is negligible.
2.5 Properties Analysis
Both implementations demonstrate unique characteristics in achieving adversarial objectives. SEED-S, by directly modifying the final reasoning step, alters only a small portion of the input, resulting in weaker attack capabilities but making it less detectable. In contrast, SEED-P demonstrates stronger attack effectiveness by generating a fully modified problem with coherent reasoning chains, showcasing broader applicability across various problem types, including multiple-choice questions and open-ended reasoning tasks.
3 Experiments
3.1 Experimental Setup
Dataset. Building on prior studies targeting reasoning processes in LLMs Xu et al. (2024); Xiang et al. (2024), we evaluate our method using four datasets that encompass diverse and challenging reasoning tasks presented in two formats. Specifically, MATH Hendrycks et al. and GSM8K Cobbe et al. (2021) focus on arithmetic reasoning with open-ended formats, while MATHQA Amini et al. (2019) presents math problems in a multiple-choice format. CSQA Talmor et al. (2019), on the other hand, features multiple-choice commonsense reasoning tasks. As for the budget constraints, we follow the approach of Xiang et al. (2024), randomly sampling 500 questions from each dataset for our experiments. Further details about datasets are provided in Appendix B.
LLMs. We evaluate four cutting-edge LLMs, encompassing both open-source and proprietary models: Llama3-8B Dubey et al. (2024), Qwen-2.5-7B Hui et al. (2024), Mistral-v0.3-7B Jiang et al. (2023), and GPT4-o Achiam et al. (2023). These models are chosen for their state-of-the-art performance and strong capabilities in solving complex reasoning tasks, providing a comprehensive benchmark to evaluate the effectiveness and versatility of our proposed attack methodology.
Settings. To assess the generalizability of SEED attack, we test its performance in both zero-shot and few-shot settings, following the traditional prompt-based Chain-of-Thought (CoT) paradigm Wei et al. (2022); Kojima et al. (2022). In the main experiments, we set $\sigma$ to 0.6, and the impact of varying $\sigma$ is explored in Section 3.3. Our experiments ’ technical specifications and implementation details are available in Appendix D.
Metrics. We assess the performance using three key metrics: accuracy (ACC), attack success rate (ASR) and detection rate. ACC measures the percentage of problems solved correctly by the model. ASR quantifies the proportion of originally correct answers that are rendered incorrect due to the attack, serving as a direct indicator of the attack’s effectiveness in disrupting the model’s reasoning capabilities. The Detection rate measures the proportion of solutions identified as originating from attacked input prompts. Further information on the metrics is available in the Appendix A.
Baselines. To the best of our knowledge, the only comparable methods are UPA and MPA proposed by Xu et al. (2024). Both approaches instruct the LLM to provide the answer first, followed by the reasoning steps. MPA goes a step further by introducing an LLM-generated false answer to mislead the model into producing corresponding incorrect reasonings. Although BadChain Xiang et al. (2024) makes a notable contribution by achieving an ASR close to 100% across all datasets, its applicability is restricted to the few-shot setting. Furthermore, as shown in Table 1, its detection ratio approaches 100%, as it simply modifies the final answer. Due to this limitation, BadChain is excluded from subsequent discussions.
Table 2: Comparison of performance measured by ASR under the setting in Xu et al. (2024). UPA and MPA are the methods proposed by Xu et al. (2024). Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Highest ASR are highlighted within each model for a given dataset setting.
Method MATH GSM8K CSQA MATHQA
Llama Z_S UPA 0.568 0.634 0.223 0.531
MPA 0.538 0.586 0.545 0.542
SEED 0.591 0.635 0.666 0.606
F_S UPA 0.682 0.719 0.107 0.570
MPA 0.674 0.653 0.400 0.689
SEED 0.732 0.745 0.572 0.718
Qwen Z_S UPA 0.418 0.414 0.210 0.527
MPA 0.437 0.486 0.308 0.545
SEED 0.473 0.495 0.324 0.511
F_S UPA 0.571 0.529 0.054 0.520
MPA 0.548 0.505 0.154 0.501
SEED 0.603 0.547 0.220 0.512
Mistral Z_S UPA 0.783 0.912 0.393 0.851
MPA 0.726 0.845 0.540 0.823
SEED 0.770 0.865 0.803 0.859
F_S UPA 0.781 0.889 0.275 0.683
MPA 0.744 0.825 0.446 0.787
SEED 0.811 0.915 0.819 0.883
GPT4-o Z_S UPA 0.249 0.212 0.109 0.473
MPA 0.307 0.237 0.177 0.447
SEED 0.326 0.295 0.512 0.482
F_S UPA 0.288 0.200 0.092 0.523
MPA 0.420 0.300 0.151 0.496
SEED 0.453 0.311 0.409 0.553
a) Performance on MATH dataset
b) Performance on CSQA dataset
Figure 4: Attack performance of SEED-P under different $\sigma$. Performance varies across models and tasks, with a range of 0.4 to 0.8 often yielding optimal results. Both lower and higher $\sigma$ values could lead to reduced ASR.
3.2 Overall Performance
Evaluation on Covert Detection As outlined in Section 1, we hypothesize that modifying $I_{solve}$ will lead to LLM outputs that are more easily identifiable, thereby diminishing the covert nature of the attack. To test this hypothesis, we evaluated the detection rates of SEED-generated solutions against baseline methods using GPT4-o. Specifically, GPT4-o was employed to classify whether the input prompt had been attacked by analyzing the corresponding output solutions. Table 1 presents detection rates on the MATH and GSM8K datasets, demonstrating that UPA and MPA fail to satisfy the covert requirements. In contrast, both SEED-S and SEED-P exhibit substantial improvements in detection rates, particularly with GPT4-o, where the average improvement exceeds 90%. As a result, GPT4-o is substantially less likely to detect the outputs of attacked LLMs under SEED attack, validating that SEED achieves a higher level of stealth by maintaining a natural reasoning flow without apparent manipulations.
Performance Comparison in Baseline Settings To ensure a fair evaluation of effectiveness, we adapted the SEED attack to match the same settings as UPA and MPA, incorporating instructions for the LLM. As shown in Table 2, SEED attack achieves improved attack performance in most cases, compared to UPA and MPA. The performance gap on CSQA is especially evident. On GPT-4o, SEED achieved an ASR more than 2x that of the baseline. This is due to the inclusion of additional reasoning steps in SEED attack that further enhance attack performance compared to UPA and MPA in most cases, indicating that SEED attack is compatible with UPA or MPA. This improvement is attributed to the SEED attack’s ability to introduce additional reasoning steps. Furthermore, these results demonstrate that SEED attack is not only a standalone approach but also compatible with methods like UPA and MPA, potentially offering a hybrid strategy for further enhancing attack performance.
Table 3: Performance comparison of the two SEED attack variations, evaluated using ACC (Accuracy) and ASR (Attack Success Rate). SEED-S and SEED-P denote SEED attack implemented through step modification and problem modification, respectively. Lower ACC and higher ASR indicate a greater impact of SEED attack. Method N represents the raw performance without any attack. Lowest ACC and highest ASR are highlighted.
Setting Method MATH GSM8K CSQA MATHQA
ACC ASR ACC ASR ACC ASR ACC ASR
Llama3 Zero_Shot N 0.541 - 0.791 - 0.680 - 0.599 -
SEED-S 0.406 0.360 0.622 0.275 0.590 0.223 0.474 0.333
SEED-P 0.370 0.514 0.520 0.425 0.302 0.626 0.382 0.518
Few_Shot N 0.528 - 0.790 - 0.710 - 0.572 -
SEED-S 0.376 0.320 0.552 0.352 0.646 0.172 0.540 0.262
SEED-P 0.374 0.496 0.444 0.503 0.394 0.516 0.360 0.531
Qwen Zero_Shot N 0.894 - 0.881 - 0.802 - 0.873 -
SEED-S 0.646 0.286 0.676 0.237 0.758 0.101 0.730 0.055
SEED-P 0.474 0.447 0.509 0.418 0.464 0.384 0.346 0.346
Few_Shot N 0.886 - 0.879 - 0.764 - 0.884 -
SEED-S 0.533 0.406 0.613 0.322 0.754 0.055 0.834 0.199
SEED-P 0.441 0.517 0.516 0.443 0.600 0.312 0.628 0.305
Mistral Zero_Shot N 0.339 - 0.520 - 0.618 - 0.403 -
SEED-S 0.223 0.500 0.180 0.672 0.506 0.251 0.190 0.670
SEED-P 0.138 0.722 0.084 0.804 0.130 0.767 0.122 0.759
Few_Shot N 0.340 - 0.468 - 0.610 - 0.366 -
SEED-S 0.231 0.563 0.296 0.543 0.566 0.210 0.334 0.536
SEED-P 0.144 0.738 0.140 0.810 0.202 0.784 0.136 0.693
GPT-4o Zero_Shot N 0.852 - 0.930 - 0.734 - 0.896 -
SEED-S 0.706 0.215 0.784 0.172 0.708 0.081 0.572 0.372
SEED-P 0.644 0.286 0.774 0.191 0.354 0.605 0.452 0.450
Few_Shot N 0.884 - 0.922 - 0.782 - 0.889 -
SEED-S 0.646 0.292 0.806 0.161 0.764 0.069 0.846 0.064
SEED-P 0.608 0.330 0.736 0.229 0.484 0.471 0.578 0.342
Effectiveness Evaluation We evaluated the effectiveness of the two implementations on all datasets and all models. As shown in Table 3, although the results vary significantly across different datasets and models, all LLMs demonstrate vulnerability under the proposed SEED attack, substantially reducing the original ACC across all tasks under both Zero-shot and Few-shot settings.
SEED-S occasionally fails, as demonstrated in CSQA and MATHQA, where the ASR is only 0.069 and 0.064 under few-shot setting, respectively. In contrast, SEED-P consistently outperforms SEED-S across all tasks. Specifically, in the CSQA and MATHQA multiple-choice datasets, SEED-S shows a low attack success rate, while SEED-P, with its more effective design, substantially boosts the ASR and reduces ACC.
Comparing the performance across different models, we find that Qwen and GPT-4o show greater robustness to the SEED attack than other models, particularly GPT-4o on MATH and GSM8K, and Qwen on CSQA and MATHQA, with ASR all under 0.4. Additionally, these models exhibit relatively higher original accuracy on the corresponding datasets. Based on these observations, we infer that a model’s performance and robustness on a task are positively correlated.
To validate this inference, we further conduct SEED-P on questions that are answered correctly and incorrectly independently. The results shown in Table 4 reveal a significant ASR gap between the two groups. The largest gap is observed in Llama-3 under the few-shot setting, achieving an ASR of 0.417. This suggests that LLMs are more robust on questions that they are capable of answering correctly, which aligns with our intuition. We also conducted experiments to investigate the transferability of SEED-P; due to space constraints, the results are presented in Appendix E.
3.3 Parameter Analysis
In the SEED attack, $\sigma$ is the hyperparameter that controls the proportion of injected reasoning steps, which intuitively influences the attack performance. To explore its impact, we evaluate the performance under different values of $\sigma$. The results are shown in Figure 4. We observe that the performance varies across different models and tasks. In most cases, a range between 0.4 and 0.6 yields competitive performance. When $\sigma$ is lower, fewer reasoning steps are injected. Consequently, the target LLM pays less attention to the injected steps and is more inclined to rely on its original reasoning process, resulting in a significant drop in ASR.
On the other hand, when $\sigma$ is higher, the ASR also drops noticeably in some cases, particularly with GPT4-o and Qwen-2.5 on MATH. We hypothesize that the reason behind this behavior is that over-injecting reasoning steps sometimes makes the LLM more robust. When too many incorrect steps are injected, the model may become aware of the inconsistencies and try to correct them, leading to a more cautious approach. This increased awareness might cause the LLM to rely more on its internal reasoning, which could prevent the attack from being as effective. More results are provided in Appendix C due to space limitation.
3.4 Ablation Study
Two key components of SEED-S are the prepending of a wrong answer and the 2-stage reasoning step generation, which involves: 1) solving the raw problem to generate the correct solution, and 2) in multiple-choice tasks, selecting a different answer and generating a corresponding solution with reasoning steps that lead to the selected answer. For open-ended tasks, the solution is directly created with reasoning steps that lead to the incorrect answer, without the need to choose a different answer.
Figure 5 illustrates the impact of these components, showing that both contribute to the overall performance. Notably, on CSQA, the 2-stage generation has a more significant effect, as in multiple-choice tasks, the LLM tends to notice when the final answer is not among the provided answer choices, prompting it to correct the error. The 2-stage reasoning generation ensures alignment between the given answer choice and the generated solution, specifically in multiple-choice tasks. More results are presented in Appendix C.
Table 4: ASR of SEED-P on questions answered correctly and incorrectly without the attack. Raw_C represents the attack performance on correctly answered questions, while Raw_I denotes the performance on incorrectly answered questions. Highest ASR are highlighted within each model for a given dataset setting.
MATH CSQA
Setting Raw_C Raw_I Raw_C Raw_I
Llama Zero_Shot 0.514 0.908 0.626 0.759
Few_Shot 0.496 0.913 0.516 0.662
Qwen Zero_Shot 0.447 0.650 0.384 0.406
Few_Shot 0.517 0.772 0.312 0.587
Mistral Zero_Shot 0.722 0.930 0.767 0.794
Few_Shot 0.738 0.942 0.455 0.823
GPT4-o Zero_Shot 0.286 0.641 0.605 0.715
Few_Shot 0.330 0.694 0.471 0.676
4 Related Work
4.1 Prompt-based Reasoning of LLMs
Enhancing the reasoning capabilities of large language models (LLMs) remains a prominent research focus Yang et al. (2024); Ning et al. . Among various approaches, the Chain of Thought (CoT) paradigm has proven particularly effective, as shown by foundational works like Wei et al. (2022) and Kojima et al. (2022). These studies demonstrated that integrating explicit reasoning steps—via reasoning exemplars or step-by-step instructions—significantly boosts LLM reasoning performance. Subsequent research has refined CoT with methods like the self-consistency strategy Wang et al. , which employs majority voting across reasoning paths, and Least-to-Most Zhou et al. , a two-stage approach that decomposes problems before solving sub-tasks. Moreover, techniques like Automatic and semi-automatic Prompting enhance CoT by optimizing prompts Zhang et al. ; KaShun et al. ; Pitis et al. (2023). Extensions to more complex structures, such as trees Yao et al. (2024) and graphs Besta et al. (2024), further expand CoT’s potential.
4.2 Prompt-based Attack on LLMs
A key area of research aimed at ensuring the safety and robustness of LLMs involves developing methods to attack these models, prompting the generation of undesirable content Deng et al. (2023); Chu et al. (2024); Yu et al. (2024). One prominent category within this field focuses on "jailbreak" attacks, which bypass alignment mechanisms to elicit harmful or unsafe outputs Yi et al. (2024); Mehrotra et al. (2023); Zheng et al. (2024). However, our work is not directly related to jailbreak attacks. Instead, we focus on adversarial attacks, which subtly manipulate outputs without noticeable input modifications Xu et al. (2022); Kandpal et al. ; Xu et al. . While earlier studies targeted traditional NLP tasks such as sentiment analysis and classification Wang et al. (2024a); Zhao et al. (2023), recent efforts have increasingly focused on attacking LLM reasoning processes Xiang et al. (2024); Xu et al. (2024). BadChain leverages backdoor vulnerabilities by embedding triggers within in-context learning demonstrations, but its applicability remains limited to specific contexts Xiang et al. (2024). Moreover, a critical drawback of BadChain is its nearly 100% detection rate, rendering it unsuitable for practical deployment. Similarly, UPA and MPA methods proposed by Xu et al. (2024), which instruct LLMs to generate answers before reasoning, often yield outputs that are easily identifiable, compromising their covert nature.
Figure 5: Ablation study of SEED-P on CSQA dataset. w/o WA: without wrong answer, w/o 2-stage: without 2-stage reasoning generation. Results show both components are important, especially 2-stage generation.
5 Conclusion and Future Works
We propose Stepwise Reasoning Error Disruption attack (SEED), a novel method targeting LLMs’ reasoning capabilities by injecting misleading steps with deliberate errors to disrupt their reasoning process. Through experiments on four datasets and LLMs, we demonstrate our method’s effectiveness with two variations, achieving high success rates while remaining stealthy. Our attack reveals LLMs’ vulnerability to adversarial reasoning steps, especially in multi-turn scenarios where early errors can cascade through the reasoning chain. Our findings highlight the need for robust defenses to protect LLMs’ reasoning integrity in critical applications.
6 Limitation
We believe our primary limitation lies in the inability to extend experiments to the entire dataset due to budget constraints. While we consider SEED to be stable and effective across various tasks, resource limitations have restricted the breadth and depth of our evaluations. Comprehensive testing across diverse datasets and scenarios would provide stronger evidence of SEED’s robustness and generalizability, which remains as our future work.
Additionally, our attack method may inadvertently generate potentially harmful or offensive content in the output solutions for the modified questions. This risk arises due to the nature of adversarial attack, which alter the model’s responses in unintended ways. Without rigorous safeguards, including targeted controls and thorough examination of outputs, the potential for generating inappropriate or harmful content cannot be fully mitigated. Future efforts should focus on integrating content moderation techniques and ethical safeguards to minimize these risks while maintaining the effectiveness of the attack method.
Appendix A Details for the Metric
Accuracy. For all datasets, exact match (EM) is used to assess the accuracy of individual problems. At the dataset level, we calculate Accuracy (ACC) to represent the percentage of problems correctly solved by the model:
$ACC=\frac{\text{Number of problems answered correctly}}{\text{Total number of problems}}.$
Attack Successful Rate. The Attack Success Rate (ASR) measures the proportion of originally correct answers that become incorrect after the attack is applied:
$ASR=\frac{|C_{\text{original}}\cap W_{\text{attack}}|}{|C_{\text{original}}|},$
Here, $C_{\text{original}}$ represents the set of correctly answered questions before the attack, and $W_{\text{attack}}$denotes the set of wrongly answered questions after the attack.This metric serves as a direct and quantitative indicator of the effectiveness of the attack in compromising the model’s reasoning capabilities.
Detection Ratio. The detection rate measures the extent to which an attack is detectable, indicating the proportion of solutions that are identified as originating from attacked input prompts. A higher detection rate suggests that the attack is more apparent, meaning it has failed to remain covert and is easily recognized.
Appendix B Details for the Datasets
MATH is a dataset of 12.5K challenging competition-level mathematics problems, each accompanied by a detailed step-by-step solution. These solutions can be used to train models to generate answer derivations and explanations Hendrycks et al. . The problems are categorized into five levels corresponding to various stages of high school. In our main experiments (Sec. 4.2), we focus on 597 algebra problems from levels 1-3 in the default test set, following  Xiang et al. (2024), and evaluate a randomly selected subset of 500 problems due to budget constraints.
GSM8K is a dataset comprising 8.5K high-quality, linguistically diverse math word problems at the grade school level, authored by human problem writers  Cobbe et al. (2021). It is divided into 7.5K training problems and 1K test problems. Each problem typically requires 2 to 8 steps to solve, involving sequences of basic arithmetic operations to determine the final answer. The problems are designed to be solvable by a capable middle school student and serve as a benchmark for multi-step mathematical reasoning. We evaluate the performance of the SEED attack on 500 randomly selected problems, constrained by the expense budget.
CSQA is a dataset designed for the commonsense question answering task. It contains 12,247 questions, each with five answer choices, requiring complex semantic understanding and often relying on prior knowledge Talmor et al. (2019). For our experiments, we use the test set provided by Diao et al. (2023b), which includes 1,221 problems and we randomly sample 500 problems for evaluation.
MATHQA is a large-scale and diverse dataset comprising 37,000 English multiple-choice math word problems spanning various mathematical domains such as algebra, calculus, statistics, and geometry Amini et al. (2019). For our experiments, we randomly sample 500 problems for evaluation due to budget constraints.
Figure 6: Prompt utilized for SEED-S and SEED-P attack, and attack detection.
Appendix C More Experiment Results
Due to space constraints, additional results from the parameter analysis and ablation study are presented in Figure 10 and Figure 7.
a) ASR on MATH dataset
b) ASR on CSQA dataset
Figure 7: Ablation study of SEED-P. w/o WA: without wrong answer, w/o 2-stage: without 2-stage reasoning generation. Results show both components are important, especially 2-stage generation on CSQA.
Appendix D Implementation of SEED attack
In Figure 6, we present the prompts employed for both the attack and problem-solving across different tasks. Additionally, in Figure 9, we display the demonstrations used in Few-Shot settings for each dataset. For a fair evaluation, it is important to note that we utilized the same demonstrations as those in Xu et al. (2024).
a) Transferability performance on MATH dataset
b) Transferability performance on CSQA dataset
Figure 8: Transferability evaluation of SEED-P on the two datasets.
Appendix E Evaluation of SEED Attack Transferability
We evaluate the transferability of the SEED attack across different datasets by conducting attacks using various LLMs on a target LLM, with the results shown in Figure 8. The results reveal that the proposed SEED attack consistently achieves a high ASR across diverse assistant and target LLM combinations, highlighting its stability and effectiveness. Furthermore, Qwen and GPT-4 stand out as the most robust target LLMs, showing relatively strong resistance to attacks from different sources. On the other hand, GPT-4 exhibits the most potent attacking capability, outperforming other models against nearly all target LLMs across datasets, especially on the CSQA dataset. This dual strength underscores GPT-4’s exceptional performance in both offensive and defensive roles.
Figure 9: Few-shot demonstration utilized for SEED-S, SEED-P attack.
a) Performance on MATH dataset
b) Performance on CSQA dataset
Figure 10: Attack performance of SEED-P under different $\sigma$. Performance varies across models and tasks, with a range of 0.4 to 0.8 often yielding optimal results. Both lower and higher $\sigma$ values could lead to reduced ASR.