### v5

# Stepwise Reasoning Disruption Attack of LLMs

Abstract
Large language models (LLMs) have made remarkable strides in complex reasoning tasks, but their safety and robustness in reasoning processes remain unexplored, particularly in third-party platforms that facilitate user interactions via APIs. Existing attacks on LLM reasoning are constrained by specific settings or lack of imperceptibility, limiting their feasibility and generalizability. To address these challenges, we propose the Stepwise rEasoning Error Disruption (SEED) attack, which subtly injects errors into prior reasoning steps to mislead the model into producing incorrect subsequent reasoning and final answers. Unlike previous methods, SEED is compatible with zero-shot and few-shot settings, maintains the natural reasoning flow, and ensures covert execution without modifying the instruction. Extensive experiments on four datasets across four different models demonstrate SEED’s effectiveness, revealing the vulnerabilities of LLMs to disruptions in reasoning processes. These findings underscore the need for greater attention to the robustness of LLM reasoning to ensure safety in practical applications. Our code is available at: https://github.com/Applied-Machine-Learning-Lab/SEED-Attack
[…]
Stepwise Reasoning Disruption Attack of LLMs
Jingyu Peng‡§111 Both authors contributed equally to this research., Maolin Wang§111 Both authors contributed equally to this research., Xiangyu Zhao§ 222 Corresponding authors. , Kai Zhang, Wanyu Wang§, Pengyue Jia§, Qidong Liu§, Ruocheng Guo, Qi Liu 222 Corresponding authors. University of Science and Technology of China, § City University of Hong Kong, Independent Researcher jypeng28@mail.ustc.edu.cn, xianzhao@cityu.edu.hk {morin.wang, jia.pengyue, wanyuwang4-c}@my.cityu.edu.hk {kkzhang08, qiliuql}@ustc.edu.cn, rguo.asu@gmail.com, liuqidong@stu.xjtu.edu.cn
1 Introduction
Large language models (LLMs) have remarkably improved complex tasks by adopting various enhanced reasoning approaches Besta et al. (2024); Xu et al. (2024b); Yang et al. (2024); Yao et al. (2024); Xu et al. (2024a); Jia et al. (2024); Cheng et al. (2024). These approaches have boosted their performance and drawn attention to the trustworthiness of the reasoning processes, including faithfulness Lanham et al. (2023); Turpin et al. (2024), fairness Shaikh et al. (2023), and safety Xu et al. (2024c).
In practice, LLMs are increasingly deployed through third-party platforms that mediate user interactions via APIs, where users do not directly access the models. This setup introduces a security risk: malicious providers could manipulate reasoning or outputs—even if model outputs seem normal at first glance, resulting in incorrect reasoning and conclusions. In this work, we investigate this specific risk by focusing on how these platforms might compromise model integrity by input manipulation.
Previous work has exposed significant LLM vulnerabilities in simple tasks such as classification and generation Wang et al. (2024); Zhao et al. (2023); Xu et al. . However, their susceptibility to attacks during the complex reasoning processes—where the stakes are often higher and the consequences are more severe in some critical areas—remains largely unexplored.
Recent advances in long reasoning methods require LLMs to iteratively build upon prior steps, facilitating reflection Madaan et al. (2024); Zhao et al. (2024) or tree search Guan et al. (2025); Zhang et al. (2024) for subsequent reasoning steps. This critical dependence on step-wise reasoning introduces a new type of vulnerability in LLMs, where manipulation of initial reasoning steps can propagate errors, causing cascading failures throughout the reasoning chain.
Exploiting such vulnerability in LLMs introduces two fundamental challenges: feasibility and imperceptibility. Technically, unlike traditional adversarial attack methods, which often leverage internal information of target models such as gradients and logits, state-of-the-art LLMs are now primarily deployed as proprietary APIs Achiam et al. (2023); Team et al. (2023). Therefore, only prompt-based attacks are feasible, where adversaries have to operate through input manipulation. While existing attempts to compromise LLM reasoning Xu et al. (2024c); Xiang et al. (2024); Ni et al. (2025) have demonstrated success in specific scenarios, they still face severe limitations in practice. A key challenge in attack design is to create attacks that are imperceptible to users. While obvious manipulations, such as altering final answers or inserting irrelevant steps, are easily detected by users, modifying the reasoning process while preserving narrative coherence is far more difficult. Existing methods often struggle to balance attack effectiveness with stealth, especially in the context of complex reasoning tasks.
Among the most relevant approaches, Xiang et al. (2024) employs misleading demonstrations to induce errors in LLMs. However, these methods are limited to in-context learning scenarios, requiring demonstrations as input, which limits their generalizability to the zero-shot settings. Furthermore, their strategy introduces an additional step that modifies the final answer, making it quite easy to identify by users. Another related approach, the preemptive answer “attack” Xu et al. (2024c) alters the reasoning paradigm of the model by producing conclusions before deriving reasoning steps. Despite its novelty, this approach often generates easily identifiable outputs, reducing its imperceptibility and effectiveness in practice. These limitations are further evidenced by our experimental results in Section 3.2.
To address these two limitations, we propose the Stepwise rEasoning Error Disruption (SEED) attack. First, SEED addresses the feasibility challenge by leveraging LLMs’ reliance on step-by-step reasoning. Instead of depending on demonstrations or backpropagated gradients, SEED strategically introduces subtle errors into the early reasoning steps. This approach achieves high success rates across a wide range of scenarios without the need for task-specific training or examples, proving its effectiveness within the constraints of proprietary API-based LLM deployments in both zero-shot and few-shot settings. Second, SEED overcomes the challenge of imperceptibility by maintaining the original prompt structure while subtly manipulating the reasoning process. The carefully introduced errors seamlessly integrate into the reasoning flow, naturally propagating through the reasoning chain to produce incorrect yet plausible-looking outcomes. This ensures that the disruptions remain covert, avoiding detection while preserving the model’s perceived trustworthiness. This novel approach not only addresses the identified limitations but also introduces a fresh perspective on how reasoning vulnerabilities in LLMs can be exploited.
Our contributions can be summarized as follows:
  • We define the task of disrupting the step-by-step reasoning process of LLMs and introduce SEED, a versatile and effective attack method that is both highly efficient in execution and challenging in detection by users.
  • We demonstrate the effectiveness and stealth of SEED across four representative LLMs on four datasets with different characteristics, which include diverse and challenging reasoning tasks presented in two different formats.
  • We naturally validate the vulnerability of LLMs to adversarially injected prior reasoning steps by designing SEED, which effectively exploits these weaknesses.
Figure 1: An example demonstrating the definition of a step-by-step reasoning task for an LLMs.
2 Method
In this section, we first provide an explicit definition of attacks that target the step-by-step reasoning process of LLMs. Following that, we introduce our two implementations of the proposed SEED attack.
Figure 2: An example of SEED-S/P attacks on a math problem. The red font highlights misleading content, where subtle calculation errors are introduced while preserving reasoning coherence. SEED-S/P first generates the initial reasoning steps containing errors, after which the target LLM produces the subsequent steps. These components are seamlessly integrated to present a complete and coherent chain-of-thought reasoning process to the user.
2.1 Problem Formulation
We first present a formal definition of a step-by-step reasoning task of LLMs as shown in Figure 1. For a given problem $p$, we define the query to the LLM, denoted as $q$, as follows:
$q=[I_{solve}\,||\,D\,||\,p],$
where $D=[d_{1},\dots,d_{K}]$ and $d_{k}$ represents the $k$-th demonstration in few-shot setting. Each demonstration $d_{k}$ is structured as $[p_{k},[r_{k}^{1},\dots,r_{k}^{T}],a_{k}]$, with $r_{k}^{t}$ being the $t$-th step in the reasoning process for the problem $p_{k}$, and $a_{k}$ representing the final answer. If $K=0$, the setting is reduced to a zero-shot scenario from few-shot.
Given $q$ as input, the corresponding output $o$ of the LLM is expressed as:
$o=LLM(q)=[R\,||\,a],$
where $R_{i}=[r^{1},\dots,r^{T}]$ is the reasoning process. Attacks targeting the reasoning process of LLMs focus on altering $o$ and its corresponding $a$ by modifying $q$ into $q^{\prime}$, which can be formulated as:
$\mathop{\arg\max}\limits_{q^{\prime}}LLM_{a^{\prime}}(q^{\prime})$ (1)
[…]
where $LLM_{a^{\prime}}$ represents the probability of the output answer being equal to $a^{\prime}$ and $\text{diff}(\cdot)$ represents the difference in terms of the narrative structure and semantic similarity.
2.2 Overview of Stepwise Reasoning Error Disruption Attack
Due to certain observations (as detailed in Section 3.2), modifications to $I_{solve}$ appear to be easily detectable, which could be partially explained by the sensitivity of the model to perturbations in problem-solving inputs. Similarly, changes to $p$ seem to be detectable by prompting the LLM to repeat the problem, potentially leveraging its tendency toward consistent reasoning in generating responses. Meanwhile, modification on demonstrations is not supported under zero-shot setting. Therefore, SEED attack performs the attack by adding misleading steps $R_{att}=[r_{att}^{1},\dots,r_{att}^{T_{att}}]$ and eliciting the LLM to output the subsequent reasoning steps $R^{\prime}=[r^{\prime 1},\dots,r^{\prime T^{\prime}}]$ and the final answer $a^{\prime}$ based on $R^{\prime}$:
$o^{\prime}=R^{\prime}||a^{\prime}=LLM([I_{solve}||D||p||R_{att}]).$
Therefore, our work focuses on how to implement a $M(\cdot)$ where $R_{att}=M(p)$, that satisfy the variation of Eq. 1:
$\mathop{\arg\max}\limits_{R_{att}}\;LLM_{a^{\prime}}(I_{solve}||D||p||R_{att})$ (2)
[…]
It’s worth noting that, as we take some reasoning steps $R_{att}$ as input, we will display $[R_{att}||R^{\prime}]$ for the victim user to maintain the integrity of reasoning process. Therefore, the constraint $\text{diff}(R,R^{\prime})$ is converted to $\text{diff}(R,[R_{att}||R^{\prime}])$.
Besides, we assume that the reasoning steps are continuous, with each step depending on the previous ones. Therefore, we can get:
$\text{diff}(R,[R_{att}||R^{\prime}]\propto\text{diff}(R[:T_{att}],R_{att}),$
with the constraint that $T_{att}+T^{\prime}=T$. In practice, as the number of reasoning steps $T$ varies, we introduce $\sigma=\frac{T_{att}}{T}$ as a hyperparameter to control the $T_{att}$. To generate $R_{att}$ that both closely resembles $R[:T_{att}]$ and effectively misleads the LLM into providing an incorrect answer, we developed two LLM-based implementations.
In the next two subsections, we introduce two implementations of the SEED attack: SEED-S (Step Modification) and SEED-P (Problem Modification). SEED-S directly alters the final step of the reasoning process, whereas SEED-P modifies the problem itself to produce the desired incorrect answer.
2.3 SEED-S: SEED Attack by Step Modification
As shown in Figure 2, one intuitive and straightforward approach is to modify the final step of $R[:T_{att}]$ with the help of an assistant LLM:
$r_{mod}$ $=LLM_{assist}(I_{mod}||p||R^{\prime}[T_{att}])$ (3)
$R_{att}$ $=R[:T_{att}-1]||r_{mod},$
where $r_{mod}$ refers to the modified reasoning step and $I_{mod}$ refers to the instruction given to the LLM to modify the reasoning step in a way that leads to an incorrect answer. It is important to note that we instruct the LLM to only modify certain digits or words related to the final answer, rather than regenerate an entirely different step, ensuring that the similarity and length constraint is still met.
However, this naive implementation has a significant limitation in terms of attack effectiveness. First, it has been observed that LLMs tend to focus more on the beginning and end of the input. As a result, they are more likely to detect inconsistencies in the final steps. Additionally, altering just a single reasoning step is often insufficient to convincingly mislead the target LLM.
Table 1: A comparison of the proportions of solutions generated by BadChain Xiang et al. (2024), UPA and MPA Xu et al. (2024c), and SEED (SEED-S and SEED-P) that were detected by GPT-4o as originating from prompts containing attacks. The average improvement is determined by calculating the average decline in the detection rate of SEED compared to Xu et al. (2024c).Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Results demonstrate that SEED methods consistently achieve substantially lower detection rates across all model architectures and settings, with SEED-P showing particularly strong stealth capabilities while maintaining attack effectiveness.
MATH GSM8K
Setting BadChain UPA MPA SEED-S SEED-P Avg. Impr. BadChain UPA MPA SEED-S SEED-P Avg. Impr.
Llama Z_S 0.998 0.382 0.440 0.170 0.252 48.7% 1.000 0.442 0.526 0.088 0.204 69.8%
F_S 1.000 0.260 0.438 0.150 0.208 48.7% 0.998 0.226 0.384 0.066 0.146 65.2%
Qwen Z_S 0.998 0.336 0.325 0.053 0.077 80.3% 0.994 0.484 0.407 0.039 0.166 77.0%
F_S 0.996 0.352 0.382 0.026 0.091 84.1% 0.996 0.439 0.497 0.042 0.162 78.2%
Mistral Z_S 0.998 0.526 0.546 0.219 0.382 43.9% 1.000 0.496 0.494 0.106 0.292 59.8%
F_S 1.000 0.537 0.478 0.212 0.421 37.6% 0.996 0.468 0.408 0.150 0.334 44.7%
GPT4-o Z_S 1.000 0.439 0.353 0.032 0.052 89.4% 1.000 0.502 0.572 0.008 0.042 95.3%
F_S 0.996 0.360 0.362 0.026 0.026 92.8% 0.998 0.426 0.406 0.014 0.022 95.7%
2.4 SEED-P: SEED Attack by Problem Modification.
To solve the limitation of SEED-S due to LLMs’ heightened attention to sequence endings and potential magnitude discrepancies in final answers, we propose a more meticulously designed implementation involving modifying the raw problem, as illustrated in Figure 2. The process begins by prompting the assistant LLM to solve the original problem and obtain the raw answer. With knowledge of this answer, the LLM is more likely to generate a modified problem that is both similar to the original and aligned with its corresponding answer. The whole process can be expressed as:
$p_{mod}||R_{mod}||a_{mod}=LLM_{assist}(p,a).$
By providing more fluent reasoning steps $R_{att}=R_{mod}[:T_{att}]$, the target LLM becomes more susceptible to being misled, ultimately producing incorrect reasoning steps and an incorrect answer.
For reasoning tasks with answer choices, the LLM is first instructed to select an answer choice, and then generate a problem based on the chosen answer. This ensures that the generated question aligns with the provided answer choices, maintaining the necessary consistency for successful attack.
To further enhance the attack’s effectiveness, inspired by Xu et al. (2024c), we prepend the corresponding incorrect answer $a_{mod}$ to $R_{att}$. Finally, the modified output of the target LLM is obtained by feeding the modified problem’s incorrect answer and partial reasoning steps into it:
$q^{\prime}=LLM(I||D||p||a_{mod}||R_{att}).$
Since we prepend $a_{mod}$ to $R_{att}$, the proportion of $a_{mod}$ relative to the entire input $q^{\prime}$ is minimal, and its position is central. Thus, we assume that its impact on the length of $R^{\prime}$ and the similarity between model outputs $R^{\prime}$ and $R[T_{att}:]$ is negligible.
It’s worth noting that although SEED-P requires $LLM_{assist}$ to initially answer the question, the accuracy of the answer has limited impact on the SEED-P’s performance. For short-answer questions, SEED-P remains effective regardless of the initial answer’s accuracy, successfully introducing faulty reasoning steps across various model performance levels. For multiple-choice questions, let the accuracy of the LLM’s responses be denoted as $P$, with a total of $K$ options for each question. While we acknowledge the theoretical constraint that the attack failure probability is $(1-P)\cdot\frac{1}{K-1}$, its effect on the model’s overall attack ability is still relatively minimal.
3 Experiments
3.1 Experimental Setup
Dataset. Building on prior studies targeting reasoning processes in LLMs Xu et al. (2024c); Xiang et al. (2024), we evaluate our method using four datasets that encompass diverse and challenging reasoning tasks presented in two formats. Specifically, MATH Hendrycks et al. and GSM8K Cobbe et al. (2021) focus on arithmetic reasoning with open-ended formats, while MATHQA Amini et al. (2019) presents math problems in a multiple-choice format. CSQA Talmor et al. (2019), on the other hand, features multiple-choice commonsense reasoning tasks. As for the budget constraints, we follow the approach of Xiang et al. (2024), randomly sampling 500 questions from each dataset for our experiments. Further details about datasets are provided in Appendix A.
Backbone LLMs. We evaluate four cutting-edge LLMs, encompassing both open-source and proprietary models: Llama3-8B Dubey et al. (2024), Qwen-2.5-7B Hui et al. (2024), Mistral-v0.3-7B Jiang et al. (2023), and GPT-4o Achiam et al. (2023). These models are chosen for their state-of-the-art performance and strong capabilities in solving complex reasoning tasks, providing a comprehensive benchmark to evaluate the effectiveness and versatility of our proposed attack methodology.
Settings. To assess the generalizability of SEED attack, we test its performance in both zero-shot and few-shot settings, following the traditional prompt-based Chain-of-Thought (CoT) paradigm Wei et al. (2022); Kojima et al. (2022). In the main experiments, we set $\sigma$ to 0.6, and the impact of varying $\sigma$ is explored in Section 3.3. Our experiments’ technical specifications and implementation details are available in Appendix B.
Metrics. We assess the performance using four key metrics: accuracy (ACC), attack success rate (ASR), modification success rate (MSR) and detection rate. ACC measures the percentage of problems solved correctly by the model. ASR quantifies the proportion of originally correct answers that are rendered incorrect due to the attack, serving as a direct indicator of the attack’s effectiveness in disrupting the model’s reasoning capabilities. MSR quantifies the proportion of problems that are altered by the attack. The detection rate measures the proportion of solutions identified as originating from attacked input prompts. Further information on the metrics is available in the Appendix C.
Baselines. To our knowledge, UPA and MPA, introduced by Xu et al. (2024c), along with BadChain Xiang et al. (2024), are the only methods targeting attacks on LLM reasoning. UPA and MPA prompt the LLM to generate an answer before the reasoning steps, with MPA further introducing a false answer to mislead reasoning. While BadChain achieves an ASR close to 100% across all datasets, its effectiveness is limited to the few-shot setting. Moreover, as Table 1 shows, its detection ratio nears 100% since it only modifies the final answer, warranting its exclusion from further discussion. Additionally, we find that the “Adding Mistake” method in Lanham et al. (2023) shares similarities with SEED-S, in that it introduces misleading reasoning steps. However, the “Adding Mistake” approach primarily focuses on examining whether CoT reasoning is post-hoc, rather than attack the reasoning of LLMs. Since the task of “Adding Mistake” differs from our single-round reasoning task, we concentrate solely on comparing the effectiveness of the attack.
Table 2: Comparison of performance measured by ASR under the setting in Xu et al. (2024c). UPA and MPA are the methods proposed by Xu et al. (2024c). Z_S and F_S stands for the Zero-Shot and Few-Shot settings, respectively. Highest ASR are highlighted within each model for a given dataset setting.
Method MATH GSM8K CSQA MATHQA
Llama Z_S UPA 0.568 0.634 0.223 0.531
MPA 0.538 0.586 0.545 0.542
SEED-P 0.591 0.635 0.666 0.606
F_S UPA 0.682 0.719 0.107 0.570
MPA 0.674 0.653 0.400 0.689
SEED-P 0.732 0.745 0.572 0.718
Qwen Z_S UPA 0.418 0.414 0.210 0.527
MPA 0.437 0.486 0.308 0.545
SEED-P 0.473 0.495 0.324 0.511
F_S UPA 0.571 0.529 0.054 0.520
MPA 0.548 0.505 0.154 0.501
SEED-P 0.603 0.547 0.220 0.512
Mistral Z_S UPA 0.783 0.912 0.393 0.851
MPA 0.726 0.845 0.540 0.823
SEED-P 0.770 0.865 0.803 0.859
F_S UPA 0.781 0.889 0.275 0.683
MPA 0.744 0.825 0.446 0.787
SEED-P 0.811 0.915 0.819 0.883
GPT-4o Z_S UPA 0.249 0.212 0.109 0.473
MPA 0.307 0.237 0.177 0.447
SEED-P 0.326 0.295 0.512 0.482
F_S UPA 0.288 0.200 0.092 0.523
MPA 0.420 0.300 0.151 0.496
SEED-P 0.453 0.311 0.409 0.553
3.2 Overall Performance
Evaluation on Covert Detection
As outlined in Section 1, we hypothesize that modifying $I_{solve}$ will lead to LLM outputs that are more easily identifiable by human, thereby diminishing the covert nature of the attack. To test this hypothesis, we evaluated the detection rates of SEED-generated solutions against baseline methods using GPT-4o to simulate human detection following the LLM-as-a-judge paradigm (Gu et al., 2024). Specifically, GPT-4o was employed to classify whether the input prompt had been attacked by analyzing the corresponding output solutions. Table 1 presents detection rates on the MATH and GSM8K datasets, demonstrating that UPA and MPA fail to satisfy the covert requirements. In contrast, both SEED-S and SEED-P exhibit substantial improvements in detection rates, particularly with GPT-4o, where the average improvement exceeds 90%. This validates that SEED achieves a higher level of stealth by preserving a natural reasoning flow without apparent manipulations. Meanwhile, SEED-S achieves lower detection rates than SEED-P, indicating superior stealth capabilities, as SEED-S introduces less modification to the prior reasoning process.
To closely simulate real-world scenarios, we conducted a rigorous human evaluation study involving 20 PhD students in computer science. Using the MATH and CSQA datasets, as detailed in Appendix D, their assessments closely aligned with our findings from GPT-4o.
We also assessed the detection rate for successfully attacked solutions, with the results presented in Section I.
a) Performance on MATH dataset
b) Performance on CSQA dataset
Figure 3: Attack performance of SEED-P under different $\sigma$. Performance varies across models and tasks, with a range of 0.4 to 0.8 often yielding optimal results. Both lower and higher $\sigma$ values could lead to reduced ASR.
Performance Comparison in Baseline Settings To ensure a fair evaluation of effectiveness, we adapted the SEED-P attack to match the same settings as UPA and MPA, incorporating instructions for the LLM. As shown in Table 2, SEED-P attack achieves improved attack performance in most cases, compared to UPA and MPA. The performance gap on CSQA is especially evident. On GPT-4o, SEED-P achieved an ASR more than 2x that of the baseline. This is due to the inclusion of additional reasoning steps in SEED-P attack that further enhance attack performance compared to UPA and MPA in most cases, indicating that SEED-P attack is compatible with UPA or MPA. This improvement is attributed to the SEED-P attack’s ability to introduce additional reasoning steps. Furthermore, these results demonstrate that SEED-P attack is not only a standalone approach but also compatible with other methods like UPA and MPA, potentially offering a hybrid strategy to further enhance attack performance.
Table 3: Performance comparison of the two SEED attack variations and “Adding Mistake” in Lanham et al. (2023), evaluated using ACC (Accuracy) and ASR (Attack Success Rate). SEED-S and SEED-P denote SEED attack implemented through step modification and problem modification, respectively. Lower ACC and higher ASR indicate a greater impact of SEED attack. Method N represents the raw performance without any attack. Lowest ACC and highest ASR are highlighted.
Setting Method MATH GSM8K CSQA MATHQA
ACC ASR ACC ASR ACC ASR ACC ASR
Llama3 Zero_Shot N 0.541 - 0.791 - 0.680 - 0.599 -
Add_M 0.414 0.345 0.625 0.272 0.568 0.230 0.498 0.310
SEED-S 0.406 0.360 0.622 0.275 0.590 0.223 0.474 0.333
SEED-P 0.370 0.514 0.520 0.425 0.302 0.626 0.382 0.518
Few_Shot N 0.528 - 0.790 - 0.710 - 0.572 -
Add_M 0.382 0.305 0.562 0.344 0.650 0.158 0.538 0.266
SEED-S 0.376 0.320 0.552 0.352 0.646 0.172 0.540 0.262
SEED-P 0.374 0.496 0.444 0.503 0.394 0.516 0.360 0.531
Qwen Zero_Shot N 0.894 - 0.881 - 0.802 - 0.873 -
Add_M 0.642 0.292 0.722 0.225 0.730 0.122 0.697 0.680
SEED-S 0.646 0.286 0.676 0.237 0.758 0.101 0.730 0.055
SEED-P 0.474 0.447 0.509 0.418 0.464 0.384 0.346 0.346
Few_Shot N 0.886 - 0.879 - 0.764 - 0.884 -
Add_M 0.546 0.394 0.672 0.285 0.730 0.086 0.874 0.133
SEED-S 0.533 0.406 0.613 0.322 0.754 0.055 0.834 0.199
SEED-P 0.441 0.517 0.516 0.443 0.600 0.312 0.628 0.305
Mistral Zero_Shot N 0.339 - 0.520 - 0.618 - 0.403 -
Add_M 0.406 0.360 0.622 0.275 0.590 0.223 0.474 0.333
SEED-S 0.223 0.500 0.180 0.672 0.506 0.251 0.190 0.670
SEED-P 0.138 0.722 0.084 0.804 0.130 0.767 0.122 0.759
Few_Shot N 0.340 - 0.468 - 0.610 - 0.366 -
Add_M 0.406 0.360 0.622 0.275 0.590 0.223 0.474 0.333
SEED-S 0.231 0.563 0.296 0.543 0.566 0.210 0.334 0.536
SEED-P 0.144 0.738 0.140 0.810 0.202 0.784 0.136 0.693
GPT-4o Zero_Shot N 0.852 - 0.930 - 0.734 - 0.896 -
Add_M 0.406 0.206 0.622 0.158 0.590 0.102 0.474 0.369
SEED-S 0.706 0.215 0.784 0.172 0.708 0.081 0.572 0.372
SEED-P 0.644 0.286 0.774 0.191 0.354 0.605 0.452 0.450
Few_Shot N 0.884 - 0.922 - 0.782 - 0.889 -
Add_M 0.673 0.254 0.818 0.158 0.730 0.083 0.872 0.045
SEED-S 0.646 0.292 0.806 0.161 0.764 0.069 0.846 0.064
SEED-P 0.608 0.330 0.736 0.229 0.484 0.471 0.578 0.342
Effectiveness Evaluation
We evaluated the effectiveness of SEED implementations and the “Adding Mistake” method across various datasets and models. As shown in Table 3, although results vary, all LLMs are vulnerable to the SEED attack, significantly reducing ACC in both zero-shot and few-shot settings. SEED-S and "Adding Mistake" perform similarly, but SEED-S generally has higher attack success rates in most cases, likely due to the summarization step in “Adding Mistake” that may alert the model to inconsistencies. SEED-S occasionally fails due to its limited ability, as seen in CSQA and MATHQA, with ASR of 0.069 and 0.064 in few-shot settings. However, SEED-P consistently outperforms SEED-S across all tasks, particularly in the CSQA and MATHQA datasets, where SEED-P greatly increases ASR and reduces ACC. This improvement is due to $LLM_{assist}$’s ability to adapt to different problems and modify key elements affecting outcomes, as shown in Appendix E.
Comparing the performance across different models, we find that Qwen and GPT-4o are more robust to the SEED attack than other models, particularly GPT-4o on MATH and GSM8K, and Qwen on CSQA and MATHQA, with ASR values all under 0.4. Additionally, these models exhibit relatively higher original accuracy on the corresponding datasets, suggesting a positive correlation between a model’s performance and robustness on a task. To validate this, we applied SEED-P separately to questions the LLM answers correctly and incorrectly, then evaluated the MSR independently. Results in Table 4 show a significant MSR gap between the two groups, with the largest gap in Llama-3 under the few-shot setting, reaching an MSR of 0.417. This indicates that LLMs are more robust on questions they answer correctly, aligning with our inference. Furthermore, the transferability evaluation presented in Appendix F confirms that more powerful LLMs can achieve both a high ASR as the assistant LLM and strong robustness as the target LLM.
In Appendix G, we evaluate self-review prompts under zero-shot settings, finding only modest improvements with ASR decreasing by no more than 10%. This suggests that simple prompt-based defenses need further refinement to counter SEED attacks. We also validated the effectiveness of prepending a wrong answer and 2-stage reasoning step generation by conducting an ablation study (see Appendix H).
3.3 Parameter Analysis
In the SEED attack, $\sigma$ is the hyperparameter that controls the proportion of injected reasoning steps, which intuitively influences the attack performance. To explore its impact, we evaluated the performance of SEED-P under different values of $\sigma$. The results, shown in Figure 3, indicate that performance varies across different models and tasks. Generally, a $\sigma$ range between 0.4 and 0.6 yields competitive performance. Lower $\sigma$ values result in fewer injected reasoning steps, causing the target LLM to rely more on its original reasoning process and leading to a significant drop in ASR.
Conversely, higher $\sigma$ values also cause noticeable ASR drops in some cases, particularly with GPT-4o and Qwen-2.5 on MATH. We hypothesize that over-injecting reasoning steps can make the LLM more robust. When too many prior steps are introduced, the LLM focuses more on reviewing its prior reasoning rather than continuing with subsequent inference. This increased scrutiny helps the LLM detect inconsistencies and attempt corrections, leading to a more cautious reasoning approach and reducing the attack’s effectiveness. Additional results are provided in Appendix J due to space limitations.
Table 4: MSR of SEED-P on questions answered correctly and incorrectly without the attack. Raw_C represents the attack performance on correctly answered questions, while Raw_I denotes the performance on incorrectly answered questions. Highest MSR are highlighted within each model for a given dataset setting.
MATH CSQA
Setting Raw_C Raw_I Raw_C Raw_I
Llama Zero_Shot 0.514 0.908 0.626 0.759
Few_Shot 0.496 0.913 0.516 0.662
Qwen Zero_Shot 0.447 0.650 0.384 0.406
Few_Shot 0.517 0.772 0.312 0.587
Mistral Zero_Shot 0.722 0.930 0.767 0.794
Few_Shot 0.738 0.942 0.455 0.823
GPT-4o Zero_Shot 0.286 0.641 0.605 0.715
Few_Shot 0.330 0.694 0.471 0.676
4 Related Work
4.1 Reasoning of LLMs
Enhancing reasoning in large language models (LLMs) remains a key research focus Yang et al. (2024); Ning et al. ; Li et al. (2023); Yuan et al. (2024); Liu et al. (2024). The Chain of Thought (CoT) paradigm has been particularly effective, as shown by Wei et al. (2022) and Kojima et al. (2022), demonstrating that explicit reasoning steps, such as exemplars or step-by-step instructions, improve LLM performance. Subsequent work refined CoT with techniques like self-consistency Wang et al. , which uses majority voting across reasoning paths, and Least-to-Most Zhou et al. , a two-stage problem decomposition approach. Further extensions to trees Yao et al. (2024) and graphs Besta et al. (2024) expand CoT’s capabilities. Recent advances in long reasoning methods require LLMs to iteratively build upon prior steps, facilitating reflection Madaan et al. (2024); Zhao et al. (2024) or tree search Guan et al. (2025); Zhang et al. (2024, 2022) for subsequent reasoning steps, further expand the reasoning ability of LLMs. This reliance on step-by-step reasoning, however, raises new concerns regarding the vulnerability of LLMs.
4.2 Prompt-based Attack on LLMs
A key area of trustworthy AI research Fan et al. (2023); Chen et al. (2022); Jia et al. (2024); Xu et al. (2025) aimed at ensuring the safety and robustness of LLMs involves developing methods to attack these models, prompting the generation of undesirable content Deng et al. (2023); Chu et al. (2024); Yu et al. (2024); Zhang et al. (2021). One prominent category within this field focuses on "jailbreak" attacks, which bypass alignment mechanisms to elicit harmful or unsafe outputs Yi et al. (2024); Mehrotra et al. (2023); Zheng et al. (2024). However, our work is not directly related to jailbreak attacks. Instead, we focus on adversarial attacks, which subtly manipulate outputs without noticeable input modifications Xu et al. (2022); Kandpal et al. ; Xu et al. . While earlier studies targeted traditional NLP tasks such as sentiment analysis and classification Wang et al. (2024); Zhao et al. (2023); Zhang et al. (2019), recent efforts have increasingly focused on attacking LLM reasoning processes Xiang et al. (2024); Xu et al. (2024c). BadChain leverages backdoor vulnerabilities by embedding triggers within in-context learning demonstrations, but its applicability remains limited to specific contexts Xiang et al. (2024). Moreover, a critical drawback of BadChain is its nearly 100% detection rate, rendering it unsuitable for practical deployment. Similarly, UPA and MPA methods proposed by Xu et al. (2024c), which instruct LLMs to generate answers before reasoning, often yield outputs that are easily identifiable, compromising their covert nature. Therefore, these approaches struggle to strike an effective balance between attack potency and stealth.
5 Conclusion and Future Works
We propose Stepwise Reasoning Error Disruption attack (SEED), a novel method targeting LLMs’ reasoning capabilities by injecting misleading steps with deliberate errors to disrupt their reasoning process. Through experiments on four datasets and LLMs, we demonstrate our method’s effectiveness with two variations, achieving high success rates while remaining stealthy. Our attack reveals LLMs’ vulnerability to adversarial reasoning steps, especially in multi-step reasoning scenarios where early errors can cascade through the reasoning chain. Our findings highlight the need for more robust defenses to protect LLMs’ reasoning integrity.
6 Limitation
We believe our primary limitation lies in the inability to extend experiments to the entire dataset due to budget constraints. While we consider SEED to be stable and effective across various tasks, resource limitations have restricted the breadth and depth of our evaluations. Comprehensive testing across diverse datasets and scenarios would provide stronger evidence of SEED’s robustness and generalizability, which remains as our future work.
Additionally, our attack method may inadvertently generate potentially harmful or offensive content in the output solutions for the modified questions. This risk arises due to the nature of adversarial attack, which alter the model’s responses in unintended ways. Without rigorous safeguards, including targeted controls and thorough examination of outputs, the potential for generating inappropriate or harmful content cannot be fully mitigated. Future efforts should focus on integrating more complicated content moderation techniques and ethical safeguards to minimize these risks while maintaining the effectiveness of the attack method.
7 Acknowledge
This research was partially supported by the National Natural Science Foundation of China (Grants No.62406303), Anhui Province Science and Technology Innovation Project (202423k09020010), Research Impact Fund (No.R1015-23), Collaborative Research Fund (No.C1043-24GF),the Fundamental Research Funds for the Central Universities, Huawei (Huawei Innovation Research Program, Huawei Fellowship), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002), Ant Group (CCF-Ant Research Fund), and Kuaishou.
Appendix A Details for the Datasets
MATH is a dataset of 12.5K challenging competition-level mathematics problems, each accompanied by a detailed step-by-step solution. These solutions can be used to train models to generate answer derivations and explanations Hendrycks et al. . The problems are categorized into five levels corresponding to various stages of high school. In our main experiments (Sec. 4.2), we focus on 597 algebra problems from levels 1-3 in the default test set, following  Xiang et al. (2024), and evaluate a randomly selected subset of 500 problems due to budget constraints.
GSM8K is a dataset comprising 8.5K high-quality, linguistically diverse math word problems at the grade school level, authored by human problem writers  Cobbe et al. (2021). It is divided into 7.5K training problems and 1K test problems. Each problem typically requires 2 to 8 steps to solve, involving sequences of basic arithmetic operations to determine the final answer. The problems are designed to be solvable by a capable middle school student and serve as a benchmark for multi-step mathematical reasoning. We evaluate the performance of the SEED attack on 500 randomly selected problems, constrained by the expense budget.
CSQA is a dataset designed for the commonsense question answering task. It contains 12,247 questions, each with five answer choices, requiring complex semantic understanding and often relying on prior knowledge Talmor et al. (2019). For our experiments, we use the test set provided by Diao et al. (2023b), which includes 1,221 problems and we randomly sample 500 problems for evaluation.
MATHQA is a large-scale and diverse dataset comprising 37,000 English multiple-choice math word problems spanning various mathematical domains such as algebra, calculus, statistics, and geometry Amini et al. (2019). For our experiments, we randomly sample 500 problems for evaluation due to budget constraints.
a) ASR on MATH dataset
b) ASR on CSQA dataset
Figure 5: Ablation study of SEED-P. w/o WA: without wrong answer, w/o 2-stage: without 2-stage reasoning generation. Results show both components are important, especially 2-stage generation on CSQA.
Figure 6: The form used for human evaluation of covert of each attack approach.
Appendix B Implementation of SEED attack
In Figure 4, we present the prompts employed for both the attack and problem-solving across different tasks. Additionally, in Figure 10, we display the demonstrations used in Few-Shot settings for each dataset. For a fair evaluation, it is important to note that we utilized the same demonstrations as those in Xu et al. (2024c).
Appendix C Details for the Metric
Accuracy. For all datasets, Exact Match (EM) is used to assess the accuracy of individual problems. At the dataset level, we calculate Accuracy (ACC) to represent the percentage of problems correctly solved by the model:
$ACC=\frac{\text{Number of problems answered correctly}}{\text{Total number of problems}}.$
Attack Success Rate. The Attack Success Rate (ASR) measures the proportion of originally correct answers that become incorrect after the attack is applied:
$ASR=\frac{|C_{\text{original}}\cap W_{\text{attack}}|}{|C_{\text{original}}|},$
where $C_{\text{original}}$ represents the set of correctly answered questions before the attack, and $W_{\text{attack}}$ denotes the set of wrongly answered questions after the attack. This metric serves as a direct and quantitative indicator of the effectiveness of the attack in compromising the model’s reasoning capabilities.
Modification Successful Rate The Modification Successful (MSR) quantifies the proportion of problems that are altered by the attack:
$MSR=\frac{|W_{\text{attack}}|}{|C_{\text{original}}\cup I_{\text{original}}|},$
where $I_{\text{original}}$ represents the set of incorrectly answered questions before the attack.
Detection Ratio. The detection rate measures the extent to which an attack is detectable, indicating the proportion of solutions that are identified as originating from attacked input prompts. A higher detection rate suggests that the attack is more apparent, meaning it has failed to remain covert and is easily recognized.
Appendix D Human Evaluation on Covert Detection
We conducted a comprehensive human evaluation study with 20 PhD students in computer science, using the MATH and CSQA datasets. Each evaluator independently assessed 10 solutions for each attack scenario, along with 10 unattacked solutions, resulting in a total of 60 evaluations per dataset. To ensure a focused and consistent evaluation, we instructed each evaluator to spend no more than 10 seconds on each solution, emphasizing the coherence of the Chain of Thought (CoT) rather than the correctness of the reasoning and calculations. The evaluation form used is shown in Figure 6.
Table 5: Human evaluation on covert detection.
BadChain MPA UPA SEED-S SEED-P Pure
MATH 0.97 0.44 0.36 0.17 0.20 0.09
CSQA 0.96 0.42 0.38 0.15 0.21 0.08
The evaluators’ assessments closely aligned with the results from our GPT-4o model, as detailed in Table 5. BadChain achieved a perfect 100% detection rate, while UPA and MPA demonstrated detection rates of 37% and 43%, respectively. In contrast, SEED-S and SEED-P showed relatively low detection rates of 16% and 21%, which were close to the 9% baseline observed for unattacked solutions.
Each PhD student spent approximately half an hour on the evaluation process, for which they were compensated US\$15 per evaluation, yielding an hourly wage of at least US\$30.
Figure 7: Case study on SEED-S/P attack. The red font highlights the modified content.
Appendix E Case Study
As shown in Figure 7, in SEED-S, $LLM_{assist}$ automatically makes modifications based on different types of problems. For instance, in mathematical problems, it modifies the intermediate calculation steps, while in multiple-choice reasoning tasks, it analyzes the options with varying degrees of inclination. However, since it can only modify one step at a time, it may not always be sufficient to persuade the LLM to output the target result. In SEED-P, $LLM_{assist}$ typically adjusts numerical values in math problems, while for common-sense reasoning tasks, it automatically identifies and modifies the most influential elements, often verbs or nouns, that affect the final outcome.
Appendix F Evaluation of SEED Attack Transferability
We evaluate the transferability of the SEED attack across different datasets by conducting attacks using various LLMs on a target LLM, with the results shown in Figure 9. The results reveal that the proposed SEED attack consistently achieves a high ASR across diverse assistant and target LLM combinations, highlighting its stability and effectiveness. Furthermore, Qwen and GPT-4o stand out as the most robust target LLMs, showing relatively strong resistance to attacks from different sources. On the other hand, GPT-4o exhibits the most potent attacking capability, outperforming other models against nearly all target LLMs across datasets, especially on the CSQA dataset. This dual strength underscores GPT-4o’s exceptional performance in both offensive and defensive roles.
Table 6: Detailed results of prompt-based self-review mitigation against SEED-P attack under zero-shot setting.
MATH GSM8K CSQA MATHQA
Llama3 SEED-P 0.514 0.425 0.626 0.518
Mitigation 0.508 0.418 0.620 0.508
Qwen SEED-P 0.447 0.418 0.384 0.346
Mitigation 0.440 0.406 0.378 0.344
Mistral SEED-P 0.722 0.804 0.767 0.759
Mitigation 0.685 0.724 0.698 0.744
GPT-4o SEED-P 0.286 0.191 0.605 0.450
Mitigation 0.276 0.184 0.568 0.432
Appendix G Prompt-based mitigation
We thoroughly tested prompt-based self-review mitigation under zero-shot setting by appending “review your reasoning steps before providing final answer” to the prompt. Our detailed results shown in Table 6 reveals modest improvement, suggesting that straightforward prompt-based defenses may require enhancement to effectively counter SEED-P attack.
Appendix H Ablation Study
Two key components of SEED-P are the prepending of a wrong answer and the 2-stage reasoning step generation, which involves: 1) solving the raw problem to generate the correct solution, and 2) in multiple-choice tasks, selecting a different answer and generating a corresponding solution with reasoning steps that lead to the selected answer. For open-ended tasks, the solution is directly created with reasoning steps that lead to the incorrect answer, without the need to choose a different answer. In the absence of the two-stage process, the LLM directly modifies the question rather than first generating the correct answer and subsequently selecting an incorrect answer for reasoning.
Figure 5 illustrates the impact of these components, showing that both contribute to the overall performance. Notably, on CSQA, the 2-stage generation has a more significant effect, as in multiple-choice tasks, the LLM tends to notice when the final answer is not among the provided answer choices, prompting it to correct the error. The 2-stage reasoning generation ensures alignment between the given answer choice and the generated solution, specifically in multiple-choice tasks.
Table 7: Comparison of the detection rate in successfully attacked solutions.
MATH GSM8K
Qwen BadChain 1.000 1.000
UPA 0.600 0.628
MPA 0.741 0.724
SEED-S 0.098 0.131
SEED-P 0.216 0.344
GPT BadChain 1.000 1.000
UPA 0.858 0.86
MPA 0.756 0.844
SEED-S 0.045 0.012
SEED-P 0.148 0.144
Appendix I The comparison of detection rate on successfully attacked solutions
In Table 7, we provide a comparison of the detection rate in successfully attacked solutions. The average detection rate is significantly higher than the overall attack detection rate, which aligns with the intuition that only modified solutions have the potential to be detected (Since ASR is computed based on questions where the original answer was correct, the detection rate for failed attacks is not zero). Additionally, the comparison between SEED and the baselines is consistent with the results presented in Table 1.
Appendix J More Experiment Results
Due to space constraints, additional results from the parameter analysis are presented in Figure 8 .
a) Performance on MATH dataset
b) Performance on CSQA dataset
Figure 8: Attack performance of SEED-P under different $\sigma$. Performance varies across models and tasks, with a range of 0.4 to 0.8 often yielding optimal results. Both lower and higher $\sigma$ values could lead to reduced ASR.
a) Transferability performance on MATH dataset
b) Transferability performance on CSQA dataset
Figure 9: Transferability evaluation of SEED-P on the two datasets.
Figure 10: Few-shot demonstration utilized for SEED-S, SEED-P attack.